THE PRODUCT OF TWO CONSECUTIVE INTERGERS IS 71 MORE THAN THE SUM. FIND THE INTERGERS
Answer by mananth(16949)

let the integers be x & x+1
x(x+1)= x+x+1+71
x^2+x=2x+72
x^2-x-72=0
x^2-9x+8x-72=0
x(x-9)+8(x-9)=0
(x-9)(x+8)=0
x= 9 OR -8
the integers are 9,10 OR -9,-10